# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### 28 Feb 2026 v4

@@ 3.1 Neighborhood bias
hence
$(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}=A_{ij}A_{j\ell},$
so $\big|(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}\big|\leq 1$. Therefore,
[…]
$\leq\frac{1}{n}+\frac{1}{n(n-1)}\big(A^{(-j)}\big)^{2}_{i\ell}.$
Since $\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell}\leq n-1$, we get
$\big|M_{i\ell}-M^{(-j)}_{i\ell}\big|\leq\frac{1}{n}+\frac{n-1}{n(n-1)}=\frac{2}{n}.$
For $i=j$ or $\ell=j$, we have $M^{(-j)}_{\mathrm{embed}}(i,\ell)=M_{i\ell}$ by definition, so the difference is zero. Hence
$\|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\leq\frac{2}{n}$
for each $j$, and taking the supremum over $j$ preserves the bound. ∎
Lemma 3.2 (Stability of expected two-hop matrices under deletion).
Let $A$ and $A^{(-j)}$ be as in Lemma 3.1, and define
$M=\frac{1}{n}A^{2},\qquad M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)}=\frac{1}{n-1}(A^{(-j)})^{2},\qquad M^{(-j)\star}=\mathbb{E}[M^{(-j)}\mid P].$
Embed $M^{(-j)}$ and $M^{(-j)\star}$ into $n\times n$ matrices as in Lemma 3.1. Then, for all $n$,
$\sup_{1\leq j\leq n}\big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}\;\leq\;\|M-M^{\star}\|_{\max}+\frac{C}{n},$ (3.4)
for a universal constant $C>0$.

### 14 Mar 2026 v5

@@ 3.1 Finite-Sample Confidence Intervals via Empirical Bernstein Bounds
Let $s_{ij}^{2}$ denote the unbiased sample variance of the neighborhood observations:
$s_{ij}^{2}=\frac{1}{h_{n}-1}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(X_{k}-\tilde{P}_{ij})^{2}.$
By Theorem 11 of Maurer and Pontil (2009), for any $\delta\in(0,1)$, the one-sided empirical Bernstein inequality bounds the upper tail of the deviation with probability at least $1-\delta$:
$\mathbb{P}\!\left(U_{ij}^{(-j)}\leq\sqrt{\frac{2s_{ij}^{2}\log(2/\delta)}{h_{n}}}+\frac{7\log(2/\delta)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\geq 1-\delta.$ (3.4)
To bound the lower tail, we apply the identical theorem to the shifted random variables $Y_{k}:=1-X_{k}$. The empirical average and expectation shift by $1$, preserving the centered deviation up to a sign change: $\tilde{P}_{Y}-\mathbb{E}[\tilde{P}_{Y}\mid\mathcal{N}_{i}^{(-j)},P]=-U_{ij}^{(-j)}$. The sample variance is translation-invariant ($s_{Y}^{2}=s_{ij}^{2}$). This yields the symmetric lower tail bound:
$\mathbb{P}\!\left(-U_{ij}^{(-j)}\leq\sqrt{\frac{2s_{ij}^{2}\log(2/\delta)}{h_{n}}}+\frac{7\log(2/\delta)}{3(h_{n}-1)}\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)\geq 1-\delta.$ (3.5)
Based on Lemma 3.1, we define the variance-adaptive half-width as
$w_{ij}^{(B)}(\alpha):=\sqrt{\frac{2s_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)},$ (3.6)
and the corresponding finite-sample confidence interval for the localized average $\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$, as
$\mathrm{CI}_{ij}^{(B)}(\alpha):=\Big[\tilde{P}_{ij}-w_{ij}^{(B)}(\alpha),\;\tilde{P}_{ij}+w_{ij}^{(B)}(\alpha)\Big]\cap[0,1].$ (3.7)
Remark 3.2 (Target of the Empirical Bernstein interval).
It is important to note that $\mathrm{CI}_{ij}^{(B)}(\alpha)$ is centered at the empirical neighborhood mean $\tilde{P}_{ij}$ and therefore provides finite-sample coverage for the localized average $\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$, not for the individual probability $P_{ij}$. The distinction reflects the fact that the Empirical Bernstein inequality controls only the stochastic fluctuation component $U_{ij}^{(-j)}$ and does not address the bias term $B_{ij}^{(-j)}$.
By construction, $\mathrm{CI}_{ij}^{(B)}(\alpha)$ captures the neighborhood average $\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$ with probability at least $1-\alpha$ for any finite $n$. To upgrade this to a coverage guarantee for the true edge probability $P_{ij}$, we must bound the deterministic bias. Under our structural assumptions, the underlying graphon is piecewise Lipschitz across a finite number of latent blocks. If a node’s latent position is too close to a block boundary, its empirical neighborhood may inadvertently capture nodes from an adjacent, structurally distinct block, compromising the smoothing procedure. To avoid these boundary effects, we restrict our entrywise inference to the regular interior of the graphon. For a fixed margin $\varepsilon>0$, let $I_{\varepsilon}$ denote the subset of the latent space strictly bounded away from all block partitions (defined formally in Section 4.3). We define the corresponding “good-edge set” as $\mathcal{E}_{\varepsilon}:=\{(i,j):i\neq j,\xi_{i}\in I_{\varepsilon},\xi_{j}\in I_{\varepsilon}\}$. As we rigorously establish in Section 4, for edges $(i,j)\in\mathcal{E}_{\varepsilon}$, the deterministic bias is uniformly bounded by a rate $r_{n}=(\log n/n)^{1/4}$.
Combining the finite-sample stochastic control with our deterministic bias bounds yields marginal coverage guarantees that are uniformly valid for any edge in the good set.
Theorem 3.3 (Marginal finite-sample inference for $P_{ij}$ on the good set).
Fix $\varepsilon>0$ and $\alpha\in(0,1)$. Let $\mathrm{CI}^{(B)}_{ij}(\alpha)$ be the empirical Bernstein interval defined in (3.7). Let $\mathcal{E}_{\varepsilon}$ denote the set of good edges defined in Section 4. Then there exist constants $C,c>0$ such that, with probability at least $1-n^{-c}$, the uniform bias bound
$\sup_{(i,j)\in\mathcal{E}_{\varepsilon}}\left|P_{ij}-\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}\right|\leq C\,r_{n}$
holds, where $r_{n}=(\log n/n)^{1/4}$. Consequently, on the intersection of the bias and variance control events, we obtain
$\mathbb{P}\!\Big(P_{ij}\in\mathrm{CI}^{(B)}_{ij}(\alpha)\oplus[-Cr_{n},\;Cr_{n}]\Big)\geq 1-\alpha-n^{-c},$ (3.8)
where $\oplus$ denotes the Minkowski sum of sets ($A\oplus B=\{a+b\mid a\in A,b\in B\}$), which explicitly widens the variance-adaptive confidence interval on both sides by the deterministic bias margin.